OpenStudy (anonymous):

square root 4x+1 -5=2 solve the equation???

OpenStudy (anonymous):

square on both sides

Parth (parthkohli):

\( \color{Black}{\Rightarrow\sqrt{4x - 1} = 2 + 5 }\) \( \color{Black}{\Rightarrow \sqrt{4x - 1} = 7 \Longrightarrow (\sqrt{4x - 1}})^2 = 7^2\)

OpenStudy (lgbasallote):

\[\sqrt{4x+1} - 5 = 2\] add 5 to both sides \[\sqrt{4x + 1} = 7\] square both sides \[4x + 1 = 49\] subtract 1 from both sides \[4x = 48\] divide both sides by 4 \[x = \frac{48}{4}\] got it?

Parth (parthkohli):

So, \(4x - 1 = 49\). Get it till this point?

OpenStudy (anonymous):

I got it...thank you

OpenStudy (lgbasallote):

welcome :D

OpenStudy (anonymous):

i am assuming you would simplify??? so your answe is 12???

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

thank you

Parth (parthkohli):

You're welcome :)